OpenStudy (anonymous):

Solve for k. k minus five-twelfths equals negative four-twelfths k equals negative three-fourths k equals negative one-twelfth k equals three-fourths k equals one-twelfth

OpenStudy (anonymous):

@Chineseboy15

OpenStudy (anonymous):

k-5/12=-4/12 k=-4/12+5/12 k=1/12 The right answer choice is the last one.

OpenStudy (anonymous):

@iceicebaby